Find the equation for each of the following items below:
a. A line that passes through (6,2) and has a slope of 2.
b. A line that passes though the points (2,4) and (8,16)
c. A line that passes through (7,3) and has a slope of -1.
d. A line that passes through (1, 1) and is perpendicular to the line 3x+4Y=12.
e. A circle with a radius of 9 and a midpoint at (3,3)

Answer by EMStelley(208)

I will do part a) for you. If you have a point (x1,y1) and a slope m, the equation for the line is
(y-y1)=m(x-x1)
So here, with a point of (6,2) and a slope of 2, we have
y-2=2(x-2)
So
y-2=2x-4
and thus
y=2x-2